OpenStudy (anonymous):

2(3)^x=3^x+3 x=-1 x=0 x=1 x=6 I think A or C

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

@Voltage

OpenStudy (anonymous):

@swolesammy

OpenStudy (anonymous):

@MsCreepyPasta

OpenStudy (anonymous):

@ronlover101

OpenStudy (anonymous):

I'm sorry,idk ;( I'm only in 6th grade :'(

OpenStudy (ronlover101):

I don't know. I am not that smart!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks anyways

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

If I subtract \(3^x\) from both sides, I get: \[\huge 2 \cdot {3^x} - {3^x} = {3^x} + 3 - {3^x}\] please simplify

OpenStudy (michele_laino):

hint: we have: \[\huge 2 \cdot {3^x} - {3^x} = \left( {2 - 1} \right){3^x} = ...\]

OpenStudy (anonymous):

2=3^x

OpenStudy (michele_laino):

what is \(2-1\)?

OpenStudy (anonymous):

1

OpenStudy (michele_laino):

correct! So, left side can be rewritten as follows: \[\huge 2 \cdot {3^x} - {3^x} = \left( {2 - 1} \right){3^x} = {3^x}\]

OpenStudy (anonymous):

so x could just equal 1?

OpenStudy (michele_laino):

therefore, the equation, can be rewritten like below: \[\huge {3^x} = {3^1}\]

OpenStudy (michele_laino):

correct! The solution is \(x=1\)

OpenStudy (anonymous):

ok thanks so much once again

OpenStudy (michele_laino):

:)